The claim is that the proportion of adults who smoked a cigarette in the past week is less than 0.25​, and the sample statistics include n=1774 subjects with 426 saying that they smoked a cigarette in the past week. Find the value of the test statistic.
Answer #1

Solution :

The null and alternative hypothesis is

H0 : p = 0.25

Ha : p < 0.25

= x / n = 426 / 1774 = 0.24

1 - P0 = 0.75

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.24 - 0.25 / [(0.25 * 0.75) / 1774]

= -0.97

Test statistic = -0.97